# Knowledge Transfer from Vision Foundation Models for Efficient Training of Small Task-specific Models

### v2

@@ C.2 MViT-V2 results
Labeled images per class 50 250 1000
Finetuned VFM (DINOv2) $47.56\pm 0.02$ $54.11\pm 0.27$ $56.95\pm 0.11$
Finetuned VFM (OpenCLIP) $48.30\pm 0.59$ $54.26\pm 0.19$ $57.03\pm 0.17$
IN-Pretrain $40.46\pm 0.07$ $47.93\pm 0.15$ $52.73\pm 0.04$
CLIP-Pretrain $42.42\pm 0.05^{*}$ $48.06\pm 0.16$ $52.72\pm 0.08$
DINOv2 Task-Agn (Patch) $41.81\pm 0.16$ $48.90\pm 0.12$ $53.15\pm 0.08$
(CC3M transfer ) Task-Agn (Image) $42.55\pm 0.19$ $48.89\pm 0.09$ $53.05\pm 0.09$
Task-Oriented $47.44\pm 0.02$ $52.44\pm 0.04$ $54.93\pm 0.04$
DINOv2 Task-Agn (Patch) $45.14\pm 0.08$ $50.37\pm 0.06$ $53.72\pm 0.08$
(Places365 transfer) Task-Agn (Image) $45.73\pm 0.02$ $50.61\pm 0.05$ $53.43\pm 0.03$
Task-Oriented $49.20\pm 0.04$ $54.02\pm 0.05$ $56.10\pm 0.03$
OpenCLIP Task-Agn $43.23\pm 0.09^{*}$ $49.07\pm 0.04$ $53.04\pm 0.14$
(CC3M transfer) Task-Oriented $47.82\pm 0.02$ $52.82\pm 0.05$ $54.92\pm 0.03$
OpenCLIP Task-Agn $47.61\pm 0.05^{*}$ $51.39\pm 0.10$ $54.17\pm 0.11$
(Places365 transfer) Task-Oriented $50.14\pm 0.04$ $54.30\pm 0.02$ $56.43\pm 0.04$
Table 9: Places365 classification accuracy for MViT-V2 target model. The results marked with $*{}^{*}$ are obtained by training only the classification layer instead of the entire model in the finetuning stage. Full finetuning produced inferior results in these cases.

### v3

@@ C.2 MViT-V2 Results
Labeled images per class 50 250 1000
Finetuned VFM (DINOv2) $47.56\pm 0.02^{\hskip 3.0pt}$ $54.11\pm 0.27$ $56.95\pm 0.11$
Finetuned VFM (OpenCLIP) $48.30\pm 0.59^{\hskip 3.0pt}$ $54.26\pm 0.19$ $57.03\pm 0.17$
IN-Pretrain $40.46\pm 0.07^{\hskip 3.0pt}$ $47.93\pm 0.15$ $52.73\pm 0.04$
CLIP-Pretrain $42.42\pm 0.05^{*}$ $48.06\pm 0.16$ $52.72\pm 0.08$
DINOv2 Task-Agn (Patch) $41.81\pm 0.16^{\hskip 3.0pt}$ $48.90\pm 0.12$ $53.15\pm 0.08$
(CC3M transfer ) Task-Agn (Image) $42.55\pm 0.19^{\hskip 3.0pt}$ $48.89\pm 0.09$ $53.05\pm 0.09$
Task-Oriented $47.44\pm 0.02^{\hskip 3.0pt}$ $52.44\pm 0.04$ $54.93\pm 0.04$
DINOv2 Task-Agn (Patch) $45.14\pm 0.08^{\hskip 3.0pt}$ $50.37\pm 0.06$ $53.72\pm 0.08$
(Places365 transfer) Task-Agn (Image) $45.73\pm 0.02^{\hskip 3.0pt}$ $50.61\pm 0.05$ $53.43\pm 0.03$
Task-Oriented $49.20\pm 0.04^{\hskip 3.0pt}$ $54.02\pm 0.05$ $56.10\pm 0.03$
OpenCLIP Task-Agn $43.23\pm 0.09^{*}$ $49.07\pm 0.04$ $53.04\pm 0.14$
(CC3M transfer) Task-Oriented $47.82\pm 0.02^{\hskip 3.0pt}$ $52.82\pm 0.05$ $54.92\pm 0.03$
OpenCLIP Task-Agn $47.61\pm 0.05^{*}$ $51.39\pm 0.10$ $54.17\pm 0.11$
(Places365 transfer) Task-Oriented $50.14\pm 0.04^{\hskip 3.0pt}$ $54.30\pm 0.02$ $56.43\pm 0.04$
Table 11: Places365 classification accuracy for MViT-V2 target model. The results marked with are obtained by training only the classification layer instead of the entire model in the finetuning stage. Full finetuning produced inferior results in these cases.
C.3 Comparison with Pseudo-labeling based Semi-supervised Learning
Table 12 shows the performance of a pseudo labeling-based semi-supervised approach with multiple rounds of training. In the first round, we obtain an initial model by training it with labeled data. In the subsequent rounds, we use the current model to generate pseudo labels for the unlabeled data and train a new model using both labeled and pseudo labeled data. In each round, the model is initialized with ImageNet pretrained checkpoint in the case of EuroSAT dataset and DINO(CC3M)-pretrained checkpoint in the case of HAM10K dataset. We use ImageNet pretraining for EuroSAT and DINO pretraining for HAM10K since these are the best performing baseline pretraining strategies for these datasets. In the case of EuroSAT classification, the performance increases when pseudo labeled data is used for the first time (round 2). However, the improvement is not significant in the next round. The final performance (87.8%) after three rounds of training is still significantly lower than 91% achieved by the proposed approach. In the case of HAM10K dataset, using pseudo labeled data results in negligible performance gain, and the proposed approach performs significantly better.
Dataset Labeled images per class Semi-supervised learning Proposed
Round 1 Round 2 Round 3
EuroSAT 5 81.4 87.2 87.8 91.0
HAM10K 500 81.6 82.4 - 86.3
Table 12: Comparison with pseudo labeling-based semi-supervised learning using FastViT target model.
C.4 Performance with Large Amount of Labeled Training Data
The main focus of this work is on improving the target task performance with limited amount of labeled data. Hence, most of our experiments focus on limited labeled data settings. In this section, we compare the performance of different approaches when large amount of target task labeled data is available.
Table 13 shows the performance achieved by various approaches on ImageNet and Places365 classification tasks when the full target task training dataset is used for both knowledge transfer and supervised finetuning (1.28M images in the case of ImageNet and 1.8M images in the case of Places365). We use FastViT as the target model for these experiments. The proposed task-oriented knowledge transfer approach outperforms all the other pretraining strategies by a significant margin in the case of ImageNet classification, and outperforms ImageNet/CLIP pretraining approaches by a significant margin in the case of Places365 classification.
Dataset VFM From scratch IN-Pretrain CLIP-Pretrain DINO-Pretrain Task-Agn Task-Oriented
ImageNet OpenCLIP 79.45 - 79.35 80.04 79.70 81.50
Places365 DINOv2 - 56.74 56.59 57.15 57.51 57.54
Table 13: Performance of FastViT target model trained with various approaches under 100% labeled training data setting.